Hayhayz (hayhayz):

Wendell is looking over some data regarding the strength, measured in Pascals (Pa), of some building materials and how the strength relates to the length. The data are represented by the exponential function f(x) = 2x, where x is the length. Explain how he can convert this equation to a logarithmic function when strength is 8 Pascals

OpenStudy (anonymous):

8 = 2^x log 8 = log 2^x = xlog 2 log 2^3 = 3log 2 = xlog 2 x = 3

Hayhayz (hayhayz):

Can you explain please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

exponential equation: nb=a logarithmic equation: logna=b

Hayhayz (hayhayz):

@jonasjl

OpenStudy (anonymous):

If you put in 8 for f(x) your exponential equation is 2x=8 Use the pattern above to change it to a log equation

Hayhayz (hayhayz):

@imqwerty

imqwerty (imqwerty):

hey wait is it \(f(x) = 2x\) or is it this- \(f(x)=2^x\)

Hayhayz (hayhayz):

The second one

imqwerty (imqwerty):

\(f(x)=2^x\) ^this is the given function and it is an exponential function and here \(f(x)\) represents the strength to make it a logarithmic function we just take "log" of both sides so we have this- \(log\left[f(x) \right] = log\left[2^x\right]\) further we can write simplify it like this- \(log\left[f(x) \right] = xlog\left[2\right]\) \(\large \frac{log\left[f(x) \right]}{log\left[2\right]} = x\) \(log_2\left[f(x) \right] = x\) this function is now a logarithmic function now its given that the strength \(\left[ f(x) \right]\) is 8 so we can write this- \(log_2(8)=x\) you can simplify this to get x

Hayhayz (hayhayz):

x=3

imqwerty (imqwerty):

yes correct :)

Hayhayz (hayhayz):

Thankkk youuu i have 4 other questions ill tag you

imqwerty (imqwerty):

okay :)